# The Six-Vertex Yang-Baxter Groupoid

### v3

@@ Abstract.
There are two main regimes of the six-vertex model: the free-fermionic point, and everything else. For the free-fermionic point, there exists a parametrized Yang-Baxter equation with a large parameter group $\operatorname{GL}(2)\times\operatorname{GL}(1)$. For non-free-fermionic six-vertex matrices, there are also well-known (group) parametrized Yang-Baxter equations, but these do not account for all possible interactions. Instead we will construct a groupoid parametrized Yang-Baxter equation that accounts for essentially all possible Yang-Baxter equations in the six-vertex model. We will also exhibit a separate groupoid for the five-vertex model. We will show how to construct solvable lattice models based on groupoid parametrized Yang-Baxter equations.
@@ 1. Introduction
Parametrized Yang-Baxter equations are a mainstay of integrable lattice models, originating in Baxter’s method of proving the commutability of row transfer matrices in the six- and eight-vertex models.
@@
In this paper we will give two groupoid parametrized Yang-Baxter equations, related to the (non-free-fermionic) six-vertex model, and the five-vertex models. These groupoids are not disjoint unions of groups.
Both of these groupoids contain well-known subgroups that are the basis of interesting Yang-Baxter equations with applications to solvable lattice models. So to demonstrate that groupoid parametrized are potentially important, in Section 8 we will consider some solvable lattice models based on an arbitrary groupoid parametrized Yang-Baxter equations. We will see that such lattice models can be constructed with both row and column solvability. These are similar to well-known models (for example [1, 10, 4]) which indeed correspond to subgroups of the groupoid. But if we work with a groupoid parametrized Yang-Baxter equation in which the groupoid is not a disjoint union of groups, the examples in Section 8 are more general.
@@ 1.1. The Six-Vertex Model
Furthermore we require that $c_{1}(u),c_{2}(u)$ is nonzero. The entries in the matrix maeq:deltatry be referred to as Boltzmann weights due to the origin of the subject in statistical mechanics [1].
@@
$\begin{array}[]{|l|l|l|l|l|l|}\hline\cr a_{1}(R)&a_{2}(R)&b_{1}(R)&b_{2}(R)&c_{1}(R)&c_{2}(R)\\ \hline\cr q_{1}z_{1}-q_{2}z_{2}&q_{1}z_{1}-q_{2}z_{2}&q_{1}(z_{1}-z_{2})&q_{2}(z_{1}-z_{2})&z_{1}w(q_{1}-q_{2})&z_{2}w^{-1}(q_{1}-q_{2})\\ \hline\cr\end{array}$
@@
So this is an example of a parametrized Yang-Baxter equation in the six-vertex model with parameter group (×$(\mathbb{C}^{\times})^{3}$.
@@
$\begin{array}[]{|l|l|l|l|l|l|}\hline\cr a_{1}(R)&a_{2}(R)&b_{1}(R)&b_{2}(R)&c_{1}(R)&c_{2}(R)\\ \hline\cr q_{1}z_{1}-q_{2}z_{2}&q_{1}z_{2}-q_{2}z_{1}&q_{1}(z_{1}-z_{2})&q_{2}(z_{1}-z_{2})&z_{1}w(q_{1}-q_{2})&z_{2}w^{-1}(q_{1}-q_{2})\\ \hline\cr\end{array}$
@@
So this is also a parametrized Yang-Baxter equation in the six-vertex model with parameter group (×$(\mathbb{C}^{\times})^{3}$.
@@
For $r,s,t$ outside $S^{\circ}$, there may be further boundary groupoids that we do not fully analyze. Particularly, there are five-vertex matrices which have themeselves important applications. We will prove in Section 7 that the matrices with $b_{2}=0$ can be organized into an interesting groupoid. The matrices with $b_{1}=0$ similarly fit into a groupoid. These lower-dimensional groupoids are not included in $\mathfrak{G}$.
This groupoid parametrized Yang-Baxter equation has the stronger property that $g\star h$ is defined if and only if $\llbracket\pi(g),w,\pi(h)\rrbracket=0$ has a solution, and if this is the case then $w$ is a constant multiple of $\pi(g\star h)$. This strong property justifies our assertion that the groupoid parametrized Yang-Baxter equation accounts for all interactions in the six-vertex model.
@@ 5. Preparations for the Non-Free-Fermionic Groupoid
It is the intersection of two quadrics and as such it is irreducible if $\alpha$ is in general position. Similarly $R_{\alpha}$ is also irreducible. Thus the fiber $\varphi^{-1}(\alpha)\cong L_{\alpha}\times R_{\alpha}$ is irreducible for $\alpha$ in general position. Thus $X$ is fibered over (×$(\mathbb{C}^{\times})^{2}$ with generically irreducible fibers, and it is therefore irreducible. ∎
@@ 6. The Six-Vertex Yang-Baxter Groupoid
From the last three results, we see that $\mathfrak{G}_{\operatorname{nf}}$ is a groupoid. We will call the disjoint union $\mathfrak{G}_{\operatorname{ff}}\sqcup\mathfrak{G}_{\operatorname{nf}}$ the six-vertex groupoid, which accounts for all Yang-Baxter equations that we may construct from the six-vertex model.
@@ 8. Solvable Lattice Models
Before we show how that $\gamma_{ij}$ can be produced, we consider the properties they need for solvability. Then we will show a systematic procedure fore producing such $\gamma_{ij}$.
@@
Here the Yang-Baxter equation is $\llbracket\pi(\rho),\pi(\gamma_{i,1}),\pi(\gamma_{i+1,1})\rrbracket=0$, so we need $\rho\star\gamma_{i+1,1}=\gamma_{i,1}$ in $\mathfrak{G}$, or $\rho=\gamma_{i,1}\gamma_{i+1,1}^{\prime}$ in $\mathfrak{G}$. Then But we want to use the Yang-Baxter equation again so we need $\rho=\gamma_{i,2}\star\gamma_{i+1,2}^{\prime}$. Thus we need

### v4

@@ Abstract.
There are two main regimes of the six-vertex model: the free-fermionic point, and everything else. For the free-fermionic point, there exists a parametrized Yang-Baxter equation with a large parameter group $\operatorname{GL}(2)\times\operatorname{GL}(1)$. For non-free-fermionic six-vertex matrices, there are also well-known (group) parametrized Yang-Baxter equations, but these do not account for all possible interactions. Instead we will construct a groupoid parametrized Yang-Baxter equation that accounts for essentially all possible Yang-Baxter equations in the six-vertex model. We will also exhibit a separate groupoid for the five-vertex model.
We will show how to construct solvable lattice models based on groupoid parametrized Yang-Baxter equations. This includes new systems that cannot be constructed using group parametrized Yang-Baxter equation. The fact that this is possible depends on the associativity axiom for a groupoid and illustrates the utility of groupoid parametrized Yang-Baxter equations.
@@ 1. Introduction
As proof of concept, in this paper we will give two groupoid parametrized Yang-Baxter equations, related to the (non-free-fermionic) six-vertex model, and the five-vertex models. These groupoids are not disjoint unions of groups.
Both of these groupoids contain well-known subgroups that are the basis of interesting Yang-Baxter equations with applications to solvable lattice models. So to demonstrate that groupoid parametrized are potentially important, in Section 8 we will consider some solvable lattice models based on an arbitrary groupoid parametrized Yang-Baxter equations. We will see that such lattice models can be constructed with both row and column solvability. These are similar to well-known models (for example [1, 10, 4]) which indeed correspond to subgroups of the groupoid. But if we work with a groupoid parametrized Yang-Baxter equation in which the groupoid is not a disjoint union of groups, the examples in Section 8 are more general. Also Zhong [13] contains an example of stochastic Yang-Baxter equations that can profitably be analyzed from the groupoid point of view.
Let us mention three questions about this work. First, we believe that the groupoid parametrized Yang-Baxter equations in this paper are not isolated example. In fact, it may be seen that there are interesting groupoid parametrized Yang-Baxter equations within the (non-field free) eight vertex model. It is possible to believe that Yang-Baxter groupoids are abundant, though perhaps not easy to find.
The second question is whether the six-vertex groupoid parametrized Yang-Baxter equation in this paper has a more uniform construction. Our construction of the groupoid in this paper depends on delicate considerations regarding the boundary components of the groupoid. Possibly the groupoid can be constructed algebraically as an action groupoid, or by other tools such as those developed in [11]. With the groupoid constructed algebraically some difficulties could perhaps be avoided in the Yang-Baxter equation.
The third question is how these Yang-Baxter groupoids fit into the theoretical framework of quantum groups or Hopf algebroids.
@@ 1.1. The Six-Vertex Model
Furthermore we require that $c_{1}(u),c_{2}(u)$ is nonzero. The entries in the matrix (3) are called Boltzmann weights due to the origin of the subject in statistical mechanics [1].
@@
[…]
@@
So this is an example of a parametrized Yang-Baxter equation in the six-vertex model with parameter group $(\mathbb{C}^{\times})^{3}$.
@@
[…]
@@
So this is also a parametrized Yang-Baxter equation in the six-vertex model with parameter group $(\mathbb{C}^{\times})^{3}$.
@@
For $r,s,t$ outside $S^{\circ}$, there may be further boundary groupoids that we do not fully analyze. Particularly, there are five-vertex matrices which have themselves important applications. We will prove in Section 7 that the matrices with $b_{2}=0$ can be organized into an interesting groupoid. The matrices with $b_{1}=0$ similarly fit into a groupoid. These lower-dimensional groupoids are not included in $\mathfrak{G}$. We are aware of at least one example of a groupoid parametrized Yang-Baxter equation outside the six-vertex model. Hence we hope that groupoid parametrized Yang-Baxter equations will be an important phenomenon.
This groupoid parametrized Yang-Baxter equation has the stronger property that $g\star h$ is defined if and only if $\llbracket\pi(g),w,\pi(h)\rrbracket=0$ has a solution, and if this is the case then $w$ is a constant multiple of $\pi(g\star h)$. This strong property justifies our assertion that the groupoid parametrized Yang-Baxter equation accounts for essentially all interactions in the six-vertex model.
@@ 5. Preparations for the Non-Free-Fermionic Groupoid
It is the intersection of two quadrics and as such it is irreducible if $\alpha$ is in general position. Similarly $R_{\alpha}$ is also irreducible. Thus the fiber $\varphi^{-1}(\alpha)\cong L_{\alpha}\times R_{\alpha}$ is irreducible for $\alpha$ in general position. Thus $X$ is fibered over $(\mathbb{C}^{\times})^{2}$ with generically irreducible fibers, and it is therefore irreducible. ∎
@@ 6. The Six-Vertex Yang-Baxter Groupoid
From the last three results, we see that $\mathfrak{G}_{\operatorname{nf}}$ is a groupoid. We will call the disjoint union $\mathfrak{G}_{\operatorname{ff}}\sqcup\mathfrak{G}_{\operatorname{nf}}$ the six-vertex groupoid, which accounts for essentially all Yang-Baxter equations that we may construct from the six-vertex model.
@@ 8. Solvable Lattice Models
Before we show how that $\gamma_{ij}$ can be produced, we consider the properties they need for solvability. Then we will show a systematic procedure for producing such $\gamma_{ij}$.
@@
Here the Yang-Baxter equation is $\llbracket\pi(\rho),\pi(\gamma_{i,1}),\pi(\gamma_{i+1,1})\rrbracket=0$, so we need $\rho\star\gamma_{i+1,1}=\gamma_{i,1}$ in $\mathfrak{G}$, or $\rho=\gamma_{i,1}\star\gamma_{i+1,1}^{\prime}$ in $\mathfrak{G}$. Then But we want to use the Yang-Baxter equation again so we need $\rho=\gamma_{i,2}\star\gamma_{i+1,2}^{\prime}$. Thus we need